Emma makes a mixture of lattes at the coffee house. The mixture must contain 25% coffee and the rest milk. If Emma has already made eighteen cups of coffee, how much milk is needed?
Answer by Theo(13342)

18 cups of coffee represents 25% of the mixture.

if we let M represent the number of cups of the mixture, we will get the following formula:

18 = .25 * M

divide both sides of this equation by .25 to get:
18 / .25 = M
simplify to get:
72 = M

the result from this formula tells us that M = 72 which means that the number of cups of Latte that can be made is 72 cups.

since the mixture must contain 25% coffee, this means that the mixture must contain .25 * 72 = 18 cups of coffee.

it does contain 18 cups of coffee so we're good up to this point.

if 72 cups of the mixture contains 18 cups of coffee and the rest is milk, then the number of cups of milk in the mixture must be equal to 72 - 18 which is equal to 54.

the mixture contains 18 cups of coffee and 54 cups of milk.

54 cups of milk are needed.

18 cups of coffee represents 25% of the mixture.
54 cups of milk represents the rest of the mixture which is equal to 75% of the mixture because 25% plus 75% is equal to 100%.

18/72 = .25 = 25% of the mixture.
54/72 = .75 = 75% of the mixture.

your solution is that 54 cups of milk is required to be added to 18 cups of coffee in order to make 72 ounces of a latte mixture.